# Q-Embroidery: A Study of Weaving Quantum Error Correction into the Fabric of Quantum Classifiers

### 20 Feb 2024 v2

@@ 1. Introduction
Quantum computing represents a significant leap forward in computational capabilities, offering the potential to solve complex problems that are intractable for classical computers. By leveraging the principles of quantum mechanics, such as superposition, entanglement and interference, quantum computers can perform certain calculations much more efficiently than their classical counterparts. This makes them particularly useful for tasks like cryptography, material science simulations, and optimization problems, where they can potentially provide solutions exponentially faster (Montanaro, 2016; Nielsen et al., 2010).
Quantum machine learning (QML) merges quantum computing with machine learning, potentially speeding up data processing and analysis tasks. This synergy could revolutionize artificial intelligence by improving big data handling, complex computations, and the creation of innovative learning algorithms. QML promises a future where quantum algorithms surpass classical methods in many areas (Schuld et al., 2015; Li et al., 2022).
As quantum computing advances, the importance of quantum error correction becomes paramount. Qubits are highly susceptible to errors due to quantum noise, which can significantly undermine the reliability of quantum computations. Quantum error correction schemes are crucial for protecting information stored in qubits from errors, thereby ensuring the practical usability and scalability of quantum computers. These error correction methods enable the construction of fault-tolerant quantum computers, making them more resilient to errors and viable for real-world applications (Terhal, 2015; Chatterjee et al., 2023).
@@ 1.2. Contribution
This research, to the best of our knowledge, is the first study to systematically apply QECC to quantum classifiers. We explore the application of three distinct QECCs: the Steane Code (Steane, 1996), and two variations of the surface code with distances $3$ and $5$ (Krinner et al., 2022). Our investigation spans a spectrum of classifier complexities, ranging from a simpler 1-qubit classifier applied to a synthetic 2D dataset to a more intricate 2-qubit classifier designed for a synthetic 4D dataset. A key element of our research is the detailed comparison of quantum circuits’ performance pre- and post-QECC application and the computational overhead each code introduces. Additionally, we conduct experiments across three error modes: depolarizing, bit and phase flips, and a combination thereof, with varying intensities of physical noise to empirically show QECC’s substantial improvement on both 1-qubit and 2-qubit classifier performance.
The importance of this seemingly simple task cannot be overstated. Implementing QECCs in even a single qubit significantly increases the need for additional qubits and gates, leading to exponential growth in overhead as circuit complexity rises. This challenge is further compounded by the need for theoretical transversal gates, such as lattice surgery, for multi-qubit gates. By concentrating on a straightforward classifier, we aim to contain this overhead and complexity, proving that success with basic models today can facilitate future progress with more complex systems as QECC technology evolves. Our findings emphasize the importance of carefully choosing QECCs based on the specific needs and constraints of quantum tasks, beyond just theoretical preferences. We assume readers have a basic understanding of QECCs, especially the Steane and Surface codes’ mechanics and applications as discussed in the existing literature (Chatterjee and Ghosh, 2024; Steane, 1996; Krinner et al., 2022).
@@ 2.1. Quantum Error Correction
Quantum Error-Correcting Codes (QECCs) stand in stark contrast to classical error correction techniques, as they safeguard information encoded in quantum states, which are defined by unique characteristics such as superposition and entanglement (Nielsen et al., 2010). The delicate nature of quantum systems means that quantum states are highly susceptible to disruption from external noise (Preskill, 1998), a vulnerability that poses a major challenge for consistent quantum computation and data storage. QECCs address this challenge effectively (Gottesman, 1997).
By spreading quantum information across multiple qubits, QECCs enable error detection and correction without necessitating a direct measurement of the quantum state, thus remaining in compliance with the no-cloning theorem (Wootters et al., 1982). Quantum computing faces primarily bit-flip and phase-flip errors, alongside more intricate errors that combine these two (Shor, 1995). Through leveraging entangled states and collective measurements, QECCs are adept at correcting such errors, enhancing the robustness and reliability of quantum information processing (Calderbank et al., 1996).
@@ 2.2. Quantum Classifiers
Quantum classifiers utilize quantum computing principles like superposition, entanglement, and interference to offer advanced data processing and analysis capabilities for machine learning tasks. These quantum properties enable the simultaneous representation and processing of extensive data combinations, intricate data correlation capture, and enhanced classification accuracy. Classifiers are mainly divided into Variational Quantum Classifiers (VQCs), which employ parameterized quantum circuits optimized through classical feedback loops, and Quantum Kernel Methods, which project input data into a high-dimensional quantum space for analysis (Schuld et al., 2015; Li et al., 2022).
The potential applications of quantum classifiers span various fields, including drug discovery for precise molecular structure analysis, finance for portfolio optimization and fraud detection, and cybersecurity for identifying complex threats beyond classical computing’s reach. Additionally, they could improve logistics and supply chain management by more efficiently solving complex optimization problems. As quantum technology progresses, its integration could lead to breakthroughs in these areas by offering solutions currently beyond classical computational methods’ capabilities (Biamonte et al., 2017).
@@ 3.2. Quantum Classifier Implementation
We partitioned the original dataset, allocating $80\%$ for training and the remaining $20\%$ for testing, employing k-fold validation to enhance the robustness of our findings. The performance metrics presented in the following sections are based on testing accuracy. The integration of rotational gates in error correction algorithms and circuits is challenging due to their sensitivity to quantum noise and the need for precise parameter control essential for quantum operation fidelity (Jayashankar et al., 2023). This necessitates transforming parameterized circuits, dependent on rotational gate fine-tuning, into non-parameterized versions. For this transformation, we employ the ‘Greedy-PQC-Optimization’ method (Phalak et al., 2024), which is designed to systematically synthesize parameterized quantum circuits into their non-parameterized equivalents, thereby sidestepping the complexities associated with parameter tuning. The comparison between the original parameterized circuits and their synthesized non-parameterized forms, focusing on any variances in accuracy and the extent of accuracy reduction following the conversion is shown in table 1. Notably, we find that the decrease in accuracy is minimal, underscoring the effectiveness of the ‘Greedy-PQC-Optimization’ process in maintaining computational performance while enhancing error resilience. Consequently, all further experiments, data analyses, and accuracy reports in this study will be based on these synthesized, non-parameterized circuits.
@@ G None 2 8 4
In our study, we focus on three specific QECCs: the Steane code (Steane, 1996) and two variations of the surface code, characterized by distances of $3$ and $5$ (Krinner et al., 2022). On one hand, the selection of the Steane code is crucial due to its unique ability to correct both bit-flip and phase-flip errors simultaneously with a relatively simple lattice structure, making it an ideal candidate for demonstrating fault-tolerant quantum computation. On the other hand, the surface code is recognized as the most feasible QECC currently available, boasting a notably high error threshold. This high threshold makes the surface code particularly attractive for practical quantum computing applications, as it suggests a greater tolerance for errors before the integrity of quantum information is compromised. To implement these QECCs in our research, we utilized the ‘mqt-qecc’ framework (Grurl et al., 2023), which allowed us to efficiently integrate these error correction schemes with the quantum classifiers previously discussed.

### 04 Mar 2024 v3

# Q-Embroidery: A Study on Weaving Quantum Error Correction into the Fabric of Quantum Classifiers

@@ 1. Introduction
Quantum computing represents a significant leap forward in computational capabilities, offering the potential to solve complex problems that are intractable for classical computers. By leveraging the principles of quantum mechanics, such as superposition, entanglement and interference, quantum computers can perform certain calculations much more efficiently than their classical counterparts. This makes them particularly useful for tasks like cryptography, material science simulations, and optimization problems, where they can potentially provide solutions exponentially faster (montanaro2016quantum, ; nielsen2010quantum, ).
Quantum machine learning (QML) merges quantum computing with machine learning, potentially speeding up data processing and analysis tasks. This synergy could revolutionize artificial intelligence by improving big data handling, complex computations, and the creation of innovative learning algorithms. QML promises a future where quantum algorithms surpass classical methods in many areas (schuld2015introduction, ; li2022recent, ).
As quantum computing advances, the importance of quantum error correction becomes paramount. Qubits are highly susceptible to errors due to quantum noise, which can significantly undermine the reliability of quantum computations. Quantum error correction schemes are crucial for protecting information stored in qubits from errors, thereby ensuring the practical usability and scalability of quantum computers. These error correction methods enable the construction of fault-tolerant quantum computers, making them more resilient to errors and viable for real-world applications (terhal2015quantum, ).
@@ 1.2. Contribution
This research, to the best of our knowledge, is the first study to systematically apply QECC to quantum classifiers. We explore the application of three distinct QECCs: the Steane Code (steane1996multiple, ), and two variations of the surface code with distances $3$ and $5$ (krinner2022realizing, ). Our investigation spans a spectrum of classifier complexities, ranging from a simpler 1-qubit classifier applied to a synthetic 2D dataset to a more intricate 2-qubit classifier designed for a synthetic 4D dataset. A key element of our research is the detailed comparison of quantum circuits’ performance pre- and post-QECC application and the computational overhead each code introduces. Additionally, we conduct experiments across three error modes: depolarizing, bit and phase flips, and a combination thereof, with varying intensities of physical noise to empirically show QECC’s substantial improvement on both 1-qubit and 2-qubit classifier performance.
The importance of this seemingly simple task cannot be overstated. Implementing QECCs in even a single qubit significantly increases the need for additional qubits and gates, leading to exponential growth in overhead as circuit complexity rises. This challenge is further compounded by the need for theoretical transversal gates, such as lattice surgery, for multi-qubit gates. By concentrating on a straightforward classifier, we aim to contain this overhead and complexity, proving that success with basic models today can facilitate future progress with more complex systems as QECC technology evolves. Our findings emphasize the importance of carefully choosing QECCs based on the specific needs and constraints of quantum tasks, beyond just theoretical preferences. We assume readers have a basic understanding of QECCs, especially the Steane and Surface codes’ mechanics and applications as discussed in the existing literature (chatterjee2024magic, ; steane1996multiple, ; krinner2022realizing, ).
@@ 2.1. Quantum Error Correction
Quantum Error-Correcting Codes (QECCs) stand in stark contrast to classical error correction techniques, as they safeguard information encoded in quantum states, which are defined by unique characteristics such as superposition and entanglement (nielsen2010quantum, ). The delicate nature of quantum systems means that quantum states are highly susceptible to disruption from external noise (preskill1998reliable, ), a vulnerability that poses a major challenge for consistent quantum computation and data storage. QECCs address this challenge effectively (gottesman1997stabilizer, ).
By spreading quantum information across multiple qubits, QECCs enable error detection and correction without necessitating a direct measurement of the quantum state, thus remaining in compliance with the no-cloning theorem (wootters1982single, ). Quantum computing faces primarily bit-flip and phase-flip errors, alongside more intricate errors that combine these two (shor1995scheme, ). Through leveraging entangled states and collective measurements, QECCs are adept at correcting such errors, enhancing the robustness and reliability of quantum information processing (calderbank1996good, ).
@@ 2.2. Quantum Classifiers
Quantum classifiers utilize quantum computing principles like superposition, entanglement, and interference to offer advanced data processing and analysis capabilities for machine learning tasks. These quantum properties enable the simultaneous representation and processing of extensive data combinations, intricate data correlation capture, and enhanced classification accuracy. Classifiers are mainly divided into Variational Quantum Classifiers (VQCs), which employ parameterized quantum circuits optimized through classical feedback loops, and Quantum Kernel Methods, which project input data into a high-dimensional quantum space for analysis (schuld2015introduction, ; li2022recent, ).
The potential applications of quantum classifiers span various fields, including drug discovery for precise molecular structure analysis, finance for portfolio optimization and fraud detection, and cybersecurity for identifying complex threats beyond classical computing’s reach. Additionally, they could improve logistics and supply chain management by more efficiently solving complex optimization problems. As quantum technology progresses, its integration could lead to breakthroughs in these areas by offering solutions currently beyond classical computational methods’ capabilities (biamonte2017quantum, ).
@@ 3.2. Quantum Classifier Implementation
We partitioned the original dataset, allocating $80\%$ for training and the remaining $20\%$ for testing, employing k-fold validation to enhance the robustness of our findings. The performance metrics presented in the following sections are based on testing accuracy. The integration of rotational gates in error correction algorithms and circuits is challenging due to their sensitivity to quantum noise and the need for precise parameter control essential for quantum operation fidelity (jayashankar2023quantum, ). This necessitates transforming parameterized circuits, dependent on rotational gate fine-tuning, into non-parameterized versions. For this transformation, we employ the ‘Greedy-PQC-Optimization’ method (phalak2024non, ), which is designed to systematically synthesize parameterized quantum circuits into their non-parameterized equivalents, thereby sidestepping the complexities associated with parameter tuning. The comparison between the original parameterized circuits and their synthesized non-parameterized forms, focusing on any variances in accuracy and the extent of accuracy reduction following the conversion is shown in table 1. Notably, we find that the decrease in accuracy is minimal, underscoring the effectiveness of the ‘Greedy-PQC-Optimization’ process in maintaining computational performance while enhancing error resilience. Consequently, all further experiments, data analyses, and accuracy reports in this study will be based on these synthesized, non-parameterized circuits.
@@ G None 2 8 4
In our study, we focus on three specific QECCs: the Steane code (steane1996multiple, ) and two variations of the surface code, characterized by distances of $3$ and $5$ (krinner2022realizing, ). On one hand, the selection of the Steane code is crucial due to its unique ability to correct both bit-flip and phase-flip errors simultaneously with a relatively simple lattice structure, making it an ideal candidate for demonstrating fault-tolerant quantum computation. On the other hand, the surface code is recognized as the most feasible QECC currently available, boasting a notably high error threshold. This high threshold makes the surface code particularly attractive for practical quantum computing applications, as it suggests a greater tolerance for errors before the integrity of quantum information is compromised. We applied QECCs on the classifiers using the ‘MQT-QECC’ framework from the Munich Quantum Toolkit (grurl2023automatic, ) and simulated the integrated circuits with IBM Qiskit’s ‘AerSimulator’.